OpenStudy (anonymous):

What is the fourth term in the expansion of (x + 2)7?

OpenStudy (anonymous):

Do you mean (x+2)^7?

OpenStudy (anonymous):

(a+b)^n=a^n +na^n-1xb +(n-1)na^n-2 b^2/2 +(n-2)(n-1)n a^n-3 b^3/6 does this help?

OpenStudy (raden):

use 7C3 (x)^(7-3)(2)^3 = 7C3*x^4*2^3 = ...